Titan Mining Corporation has 7.1 million shares of common stock outstanding, 255,000 shares of $4.30 annual dividend preferred stock outstanding, and 140,000 bonds with a semiannual coupon rate of 5.6 percent outstanding, par value $1,000 each. The common stock currently sells for $66 per share and has a beta of 1.10, the preferred stock currently sells for $90 per share, and the bonds have 19 years to maturity and sell for 108 percent of par. The market risk premium is 7.6 percent, T-bills are yielding 2.9 percent, and the company’s tax rate is 25 percent.

a.

What is the firm’s market value capital structure? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., .1616.)

b. If the company is evaluating a new investment project that has the same risk as the firm’s typical project, what rate should the firm use to discount the project’s cash flows?
Answer #1

(a) Bond (Debt):

Par Value = $ 1000, Market Value = 108% of Par Value = 1.08 x 1000 = $ 1080, Semi-Annual Coupon Rate = 5.6 % per annum,

Bond Tenure = 19 years or (19 x 2) = 38 half-years, Semi-Annual Coupon = 0.5 x 0.056 x 1000 = $ 28

Let the yield to maturity of the bond be 2y% and number of bonds issued be 140000

Therefore, 1080 = 28 x (1/y) x [1-{1/(1+y)^(38)}] + 1000 / (1+y)^(38)

Using EXCEL's Goal Seek Function to solve the above equation, we get:

y = 0.02473 or 2.473 %

Yield to Maturity = 2 x y = 2 x 2..473 = 4.946 %

Cost of Debt = kd = Yield to Maturity = kd = 4.946 %

Market Value of Debt = 140000 x 1080 = $ 151200000 or $ 151.2 million

Common Stock:

Common Stock Price = $ 66, Number of Shares Outstanding = 7.1 million

Market Value of Common Stock = 66 x 7.1 = $ 468.6 million

T-Bill Yield = Risk-Free Rate = Rf = 2.9 %, Market Risk Premium = MRP = 7.6 %, beta = 1.1

Therefore, as per CAPM = Cost of Equity = ke = Rf + Beta x MRP = 2.9 + 1.1 x 7.6 = 11.26 %

Preferred Stock:

Preferred Stock Price = $ 90, Annual Dividend = $ 4.3

Cost of Preferred Stock = kp = (4.3/90) = 0.04778 or 4.778 %

Number of Preferred Stock = 255000

Market Value of Preferred Stock = 255000 x 90 = $ 22950000 or $ 22.95 million

Market Value of Firm = 22.95 + 468.6 + 151.2 = $ 642.75 million

Market Value Capital Structure:

Common Stock Proportion = E = (468.6/642.75) = 0.729

Preferred Stock Proportion = P = (22.95/642.75) = 0.036

Debt Proportion = D = (151.2 / 642.75) = 0.235

(b) If a firm is evaluating a new investment project which has risk levels comparable to that of the regular projects and firm risks, then the appropriate discount rate for the project cash flows should be the firm's weighted average cost of capital (WACC)

Tax Rate = t = 25 %

WACC = kp x P + ke x E + kd x (1-t) x D + ke x E = 4.778 x 0.036 + 11.26 x 0.729 + 4.946 x (1-0.25) x 0.235 = 9.25228 % ~ 9.25 %